How many moles of gas exist if it has a volume of 500 mL at 23 degrees Celsius and 23 psi?
1
Expert's answer
2020-06-04T10:38:11-0400

Using Ideal Gas law, "PV=nRT \implies n=PV/RT"

"P=23psi= 23*0.068046 atm"

"V= 500 ml = 0.5 L"

"T=23^o C= 273+23=296 K"

"R= 0.0821 Latmmol^{-1}K^{-1}"

"\implies n = 23*0.068046*0.5/(0.0821*296)= 0.0322 mol"